OpenStudy (anonymous):

What is the vertex of the graph of y = |–x – 3|+ 1?

OpenStudy (jdoe0001):

what value of "x" makes -> |–x – 3| < - equals to 0?

OpenStudy (anonymous):

i really dont know.. they gave me 4 options (3, –1) (4, 0) (0, 4) (–3, 1)

OpenStudy (anonymous):

The minimum value of an absolute value function is 0. @jdoe0001 was asking what value of x gives you the minimum value of |-x - 3|. Knowing this will give you the solution you desire.

OpenStudy (jdoe0001):

well, look at the absolute value expression

OpenStudy (anonymous):

(0,4)? that would be the answer right cayse x is zero?

OpenStudy (jdoe0001):

well.... I mean. what value of "x" makes \(\huge |–x – 3| = 0\)

OpenStudy (anonymous):

i dont know thats why i need help

OpenStudy (anonymous):

Maybe if the absolute value signs were gone it would help. Solve -x - 3 = 0.

OpenStudy (jdoe0001):

yeap... so... what do you think? \(\huge -x -3 = 0\)

OpenStudy (anonymous):

so x would be 3 and y would be -1

OpenStudy (anonymous):

-x - 3 = 0. Check if x = 3 is correct.

OpenStudy (jdoe0001):

-x-3 =0 # x = 3 \(\huge -3-3\ne 0\)

OpenStudy (anonymous):

so which one is it ?????? (3, –1) (4, 0) (0, 4) (–3, 1)